### v1

# ScVLM: a Vision-Language Model for Driving Safety Critical Event Understanding

Abstract
Accurately identifying, understanding, and describing driving safety-critical events (SCEs), including crashes and near-crashes, is crucial for traffic safety, automated driving systems, and advanced driver assistance systems research and application. As SCEs are rare events, most general Vision-Language Models (VLMs) have not been trained sufficiently to link SCE videos and narratives, which could lead to hallucination and missing key safety characteristics. To tackle these challenges, we propose ScVLM, a hybrid approach that combines supervised learning and contrastive learning to improve driving video understanding and event description rationality for VLMs. The proposed approach is trained on and evaluated by more than 8,600 SCEs from the Second Strategic Highway Research Program Naturalistic Driving Study dataset, the largest publicly accessible driving dataset with videos and SCE annotations. The results demonstrate the superiority of the proposed approach in generating contextually accurate event descriptions and mitigate hallucinations from VLMs.
Keywords Driving Safety-Critical Events  $\cdot$ Vision-Language Models  $\cdot$ Supervised Learning  $\cdot$ Contrastive Learning  $\cdot$ Event Description Rationality
1 Introduction
In the domain of autonomous driving, VLMs have demonstrated strong and robust capabilities in perception, scene understanding, decision-making, and adaptability to novel scenarios [1, 2, 3, 4]. Related to driving safety, VLMs proficiently interpret environmental information surrounding the vehicle and possess foundational insights into traffic accidents and potential risk factors [2, 4]. Despite these advances, VLMs still face challenges in accurately identifying safety-critical events (SCEs), including crashes and near-crashes. Furthermore, understanding the nature of these SCEs, such as conflicts with a lead vehicle, remains elusive. This information is crucial for assessing driving safety.
Figure 1: Example scene understanding result by VideoLLaMA2
Figure 1 illustrates the capabilities of the advanced VLM, Video Large Language Model for Learning from All Modalities 2 (VideoLLAMA2) [5], in understanding SCEs. This model exhibits a excellent understanding of static environmental contexts, including weather conditions and the immediate surroundings. However, its ability to discern dynamic elements crucial for SCE analysis—such as distinguishing between crash and normal driving scenarios or identifying the nature of conflicts (e.g., with a leading or parked vehicle)—is still constrained. These findings underscore the necessity for improved models capable of more effectively interpreting dynamic information in SCE videos.
The limitations observed in VideoLLAMA2’s performance on SCE analysis can be attributed to two key factors. The rarity of SCEs in real-world scenarios results in insufficient training data for general VLMs to establish connections between SCE videos and corresponding narratives [6, 7]. The scarcity of relevant training examples can lead to hallucination and the omission of crucial safety characteristics in the model’s interpretations. Additionally, the abstract nature of event types and conflict types poses a significant challenge for VLMs to accurately identify [8].
This work introduces a novel hybrid approach that utilizes supervised learning, contrastive learning, and VLM for driving video understanding, with emphasis on SCEs. For event type identification, including crashes, tire strikes, near-crashes, and normal driving, supervised learning techniques are used due to their task-specific effectiveness. For conflict type identification, contrastive learning is employed, taking advantage of semantic label dependency on labels with rich text information. To interpret the surrounding environmental context, VLM is utilized for its capability to accurately identify concrete objects in the video. Finally, VLM combines the predictions from the supervised and contrastive learning approaches with the environmental context to generate a coherent event description in fluent language.
The primary contribution of this study is the development of an accurate event description generator that addresses the issue of hallucinations in VLMs. The proposed approach enhances prediction precision for these elements, thereby guiding the VLM to generate more accurate event descriptions.
The evaluation of the proposed approach utilized data from the Second Strategic Highway Research Program (SHRP 2) Naturalistic Driving Study (NDS), which is the largest publicly accessible NDS dataset to date, containing over 1,000,000 hours of continuous driving data [9]. The SHRP 2 NDS data includes rich driving information from multiple cameras, kinematic sensors, radar, and GPS. From the continuous driving data, a dedicated project was conducted to identify SCEs and randomly selected normal driving baselines [9], including four distinct event types: crashes, tire strikes, near-crashes, and normal driving baselines. SCEs went through a rigorous data annotation process to extract the nature of conflict. The annotation provides detailed conflict type labels for SCEs, covering scenarios like conflicts with a lead vehicle, single-vehicle conflicts, and conflicts with a vehicle turning into another’s path in the same direction. This rich dataset is ideal for evaluating the effectiveness of the proposed hybrid approach.
2 Related Works
VLM for Driving Scene Understanding VLMs combine visual and language processing to interpret driving scenarios and aid decision-making. DriveVLM, which incorporates reasoning modules for scene description and analysis, addressing spatial reasoning and computational challenges by proposing a hybrid system combining VLMs with traditional autonomous driving pipelines [3]. DriveScenify, utilizing advanced VLMs to generate contextually relevant responses based on driving scene videos, aiming to enhance urban mobility and road safety [10]. Shoman et al. [2] propose a parallel architecture that integrates object detection, tracking, and natural language generation to produce detailed descriptions of traffic events, thereby improving traffic safety through comprehensive event analysis. Jain et al. [11] integrate VLMs with multi-sensor data to enhance the comprehension of traffic dynamics and interactions among road users and infrastructure.
While there has been a significant expansion in research on VLMs for general scene understanding in driving contexts, the specific focus on SCEs, which are crucial for enhancing safety and reliability in autonomous vehicles, is relatively under-explored. Some researches that mention crashes or traffic accidents [2, 4, 3] do not explore the intricacies of these events in depth.
Supervised Learning and Contrastive Learning Supervised learning and multi-modal contrastive learning are two popular approaches for driving video scene classification tasks [12, 13, 14, 15, 16]. Supervised learning relies on one-hot or figure-coded labels to train models [17], while multi-modal contrastive learning, particularly in a video-text manner, leverages the relationships between different modals of the data to learn useful representations [18]. In supervised learning, state-of-the-art and efficient methods such as SlowFast [19], Swin Transformer [20], and TimeSformer [21] have proven effective for video scene understanding. Meanwhile, in contrastive learning, inspired by CLIP [22], notable approaches like X-CLIP [23] and ActionCLIP [24] excel in video understanding, particularly in few-shot tasks. The X-CLIP introduced a lightweight cross-frame attention mechanism and proposed a video-adaptive textual prompting scheme to handle video-text datasets [23]. The ActionCLIP introduced textual and visual adapters to enhance the model’s ability to process and understand text and video modalities [24].
3 VLM-based Driving SCE Analysis
To generate a comprehensive and accurate description of SCEs, the proposed approach is divided into three stages, as illustrated in Figure 2. The first stage employs a supervised learning approach to analyze the front-view videos and classify four types of events – crashes, tire strikes, near-crashes, and normal driving. In the second stage, a contrastive learning approach is utilized to identify 16 types of conflicts, such as conflict with a lead vehicle, a parked vehicle, or a following vehicle. The final stage integrates the event and conflict type information into a VLM to generate a comprehensive description of the event.
Figure 2: Overall Model Process
3.1 Supervised Learning for Event Type Classification
Supervised learning for event type classification from video is a 1-of-N vote problem, as illustrated in Figure 3. This type of model takes a video as input and feeds it through a video encoder to generate video representation. The representation is subsequently processed by a classifier to produce prediction scores. The model is optimized by minimizing the cross entropy loss based on the prediction scores. Formally, given an input video $x_{i}$ and a label $y_{i},i=1,2,...,N$ from a predefined set of labels $Y$, supervised learning approaches typically train a model to estimate the conditional probability $P(y|x,\theta)$.
Figure 3: Supervised learning structure for video data.
The supervised learning approach employs a video encoder $g_{V}$, which extracts representations for video data. Then the classifier projects the video representations into the space with the dimension of labels to obtain the prediction scores:
$s(x)=\text{Classifier}[g_{V}(x)]$ (1)
Subsequently, the loss to be optimized is defined as the cross-entropy loss between prediction scores and the ground truth:
$\begin{split}L=\text{Cross Entropy}[s(x),y]\end{split}$ (2)
where the ground-truth label $y$ is converted into a numerical representation or a one-hot vector that indicates its position within the entire label set of length $|Y|$. During the inference phase, the index with the highest score from the predictions is considered the corresponding category.
3.2 Contrastive Learning for Conflict Type Classification
Contrastive learning approach is illustrated in Figure 4. This approach processes a video-text pair as input. The input video is fed into the video encoder to generate video representations. Concurrently, the label text is fed into the text encoder to obtain text representations. Contrastive learning approach computes a similarity score matrix between the video and text representations and is optimized by minimizing the loss between this similarity matrix and the ground-truth video-text pair matrix.
Figure 4: Contrastive learning structure for video-text pair data.
In contrastive learning, the video classification task is redefined as predicting the probability $P[f(x,y)|\theta]$, where $y$ represents the original label texts and $f$ denotes a similarity function. Subsequently, the inference becomes a matching process, with the label texts having the highest similarity score being the outcome:
$\hat{y}=\arg\max_{y\in Y}P[f(x,y)|\theta]$ (3)
Contrastive learning approach employs separate encoders $g_{V}$ and $g_{T}$ for videos and label texts within a dual-stream framework. The video encoder $g_{V}$ extracts spatio-and-temporal representations for video data, while the language encoder $g_{T}$ captures representations from label texts. To bring matched video and label representations closer, the similarity score is defined using cosine distances:
$s(x,y)=\frac{v^{T}t}{\|v\|\|t\|},\quad s(y,x)=\frac{t^{T}v}{\|t\|\|v\|}$ (4)
where $v=g_{V}(x)$ and $t=g_{T}(y)$ represent the encoded representations of $x$ and $y$, respectively. Subsequently, the softmax-normalized video-to-text and text-to-video similarity scores are computed as:
$\begin{split}p_{x\to y_{i}}(x)=\text{SoftMax}[s(x,y_{i})]\\ p_{y\to x_{i}}(y)=\text{SoftMax}[s(y,x_{i})]\end{split}$ (5)
The ground-truth similarity scores are denoted as $q_{x\to y_{i}}(x)$ and $q_{y\to x_{i}}(y)$, respectively. The negative pair has a similarity of 0, and the positive pair has a similarity of 1. The video-text contrastive loss to be optimized is defined as
$\begin{split}L=\frac{1}{2}\mathbb{E}_{(x,y)\sim D}[l(p_{x\to y}(x),q_{x\to y}(x))\\ +l(p_{y\to x}(y),q_{y\to x}(y))]\end{split}$ (6)
where $D$ is the training set; $l$ is either cross-entropy loss (for single-label dataset) or Kullback–Leibler (KL) divergence (for multi-label dataset).
Model trained by contrastive learning approach can carry out inference, as illustrated in Figure 5. When presented with a testing dataset with a label set comprising $M$ labels, the initial step involves extracting the label representations, $[t_{k}],k=1,2,...,M$, using the language encoder, $g_{T}$. Subsequently, for a given testing video, its representation $v$ is obtained through the video encoder, $g_{V}$. The similarity between $v$ and each label representation $t_{k}$ is computed using Equation (4). The label assigned to the video is determined as the one with the highest similarity score with $v$.
Figure 5: Inference procedure of contrastive learning approach.
3.3 VLM for Event Narrative Generation
In this study, VLM is utilized to generate narrative descriptions based on contextual factors, such as weather conditions, geographical location, and surrounding environment. The process involves the VLM performing inference when provided with a text prompt and the result of SCE detection, enabling accurate event description.
Formally, the video representation $V_{i}$ for a given input video $x_{i}$ is obtained using a video encoder $g_{V}$:
$\begin{split}V_{i}=g_{V}(x_{i})\end{split}$ (7)
The visual representation $V_{i}$ is subsequently processed through the Spatial-Temporal Convolution (STC) connector to capture spatial-temporal dynamics. Given a text prompt $P$, along with the predicted event type $\hat{y_{et}}$ and conflict type $\hat{y_{ct}}$ (if applicable), the output response is generated from a pretrained large language model (LLM):
$\begin{split}R=\text{LLM}[\text{STC}(V_{i}),P,\hat{y_{et}},\hat{y_{ct}}]\end{split}$ (8)
4 Application and Results
Problem Setup Utilizing the SHRP 2 NDS dataset [9], this study focuses on generating accurate narratives from front-view video of SCEs. Normal driving segments are captured a few seconds before or after SCEs within the trip as a reference level. In total, the dataset includes 1,063 crashes, 774 tire strikes, 6,782 near-crashes, and 8,497 normal driving segments, with SCEs (crashes, tire strikes, and near-crashes) classified into 16 conflict types. Each event consists of 30-second front-view video.
The proposed approach aims to address three tasks: a classification task to distinguish event types, a classification task to differentiate conflict types, and a task to generate event text narrative. To the best of the authors’ knowledge, this is the only publicly available driving video dataset with labeled event and conflict types suitable for these three tasks, thereby supporting the evaluation of the proposed approach.
4.1 SHRP 2 NDS Dataset
Model Learning Approach Accuracy mAP AUC Balanced Accuracy Macro Precsion Macro F1
X-CLIP Contrastive 0.829 0.708 0.937 0.653 0.688 0.666
Action CLIP Contrastive 0.816 0.659 0.901 0.639 0.646 0.642
SlowFast Supervised 0.917 0.862 0.981 0.787 0.811 0.797
Swin Transformer Supervised 0.894 0.810 0.969 0.738 0.776 0.755
TimeSformer Supervised 0.851 0.727 0.950 0.650 0.691 0.668
Table 1: Comparison of different models in event type classification
Model Learning Approach Accuracy Top5 Acc mAP AUC Balanced Acc Macro Precsion Macro F1
X-CLIP Contrastive 0.766 0.951 0.547 0.921 0.493 0.599 0.508
Action CLIP Contrastive 0.748 0.949 0.488 0.907 0.439 0.520 0.458
SlowFast Supervised 0.721 0.928 0.467 0.927 0.437 0.469 0.423
Swin Transformer Supervised 0.719 0.927 0.432 0.889 0.411 0.450 0.420
TimeSformer Supervised 0.713 0.945 0.468 0.920 0.448 0.487 0.459
Table 2: Comparison of different models in conflict type classification
The SHRP 2 NDS represents the largest naturalistic driving study to date, collecting data from over 3,000 participants [9, 25]. Participants’ personal vehicles were equipped with an integrated data collection system featuring four cameras (front, driver’s face, over-the-shoulder, and rear views). Video data was continuously recorded at 15 frames per second (FPS) from vehicle start to shutdown, containing over 1,000,000 hours or 70 million miles of driving data.
ID Conflict type Count
1 Conflict with a lead vehicle 3165
2 Single vehicle conflict 1441
3 Conflict with vehicle turning into another
vehicle path (same direction) 377
4 Conflict with parked vehicle 173
5 Conflict with vehicle in adjacent lane 1508
6 Conflict with vehicle turning across another
vehicle path (opposite direction) 242
7 Conflict with a following vehicle 181
8 Conflict with vehicle turning into another
vehicle path (opposite direction) 316
9 Conflict with vehicle moving across another
vehicle path (through intersection) 170
10 Conflict with animal 360
11 Conflict with vehicle turning across another
vehicle path (same direction) 65
12 Conflict with merging vehicle 121
13 Conflict with pedal cyclist 64
14 Conflict with pedestrian 163
15 Conflict with obstacle/object in roadway 176
16 Conflict with oncoming traffic 78
17 Unknown 19
Table 3: Count of SCEs by conflict types
SCEs, including crashes, tire strikes, and near-crashes, were identified through a multi-step process. This involved evaluating the kinematic characteristics of all driving data and verifying SCEs via video analysis by trained data analysts [9]. A near-crash is defined as a situation requiring an evasive maneuver by any party to avoid a crash [9], while a tire strike event is linked to a road departure incident [26]. The 8,619 SCEs are categorized into 16 conflict types, as detailed in Table 3. After excluding the ’unknown’ type, the total is 8,600. The detailed definitions of these conflict types can be found on the SHRP 2 InSight website [27].
4.2 Application
Data Pre-processing The temporal location of each SCE is pinpointed using the impact timestamp from the SHRP 2 database and serves as the center point of the event [27]. A temporal window encompassing 38 video frames (equivalent to 2.5 seconds) both preceding and succeeding the event was extracted, resulting in a 5-second interval of the front-view video. The normal driving segment, randomly selected from either before or after the SCE, consists of 77 video frames to match the duration of the SCE segment.
Model Implementation For both classification tasks, the dataset was randomly divided into training, testing, and validation subsets in a 7:2:1 ratio. For the event type classification task, this corresponds to 11,981 events for training, 3,424 for testing, and 1,711 for validation. For the conflict type classification task, there are 6,030 SCEs for training, 1,724 for testing, and 861 for validation. The validation sets were used for hyper-parameter tuning, while evaluation performance was assessed on the independent testing sets. The software environment was based on Python 3.8 running on Rocky Linux 9.3. The model was trained on a high-performance GPU workstation equipped with dual Intel Xeon Gold 6338 CPUs @ 2.00 GHz, 256 GB RAM, and two Nvidia Tesla A100 80 GB GPUs.
This study compares the performance of supervised learning and contrastive learning approaches on two classification tasks using five benchmarks. The selected contrastive learning algorithms are X-CLIP [23] and ActionCLIP [24], while the supervised learning algorithms include SlowFast [19], Video Swin Transformer [28], and TimeSformer [29]. These methods demonstrate competitive performance on the Kinetics-400 dataset [30], providing a suitable basis for evaluating the efficacy of the two learning approaches in two classification tasks.
X-CLIP adopts the ViT-B/16 CLIP architecture as parts of the cross-frame communication transformer and a one-layer multi-frame integration transformer [23]. ActionCLIP utilizes the ViT-B/16 CLIP architecture with six Transformer adapter layers [24]. The SlowFast model employs a ResNet3D backbone structure [19]. The Video Swin Transformer is based on the Swin-Base model architecture [28]. The TimeSformer employs a TimeSformer-Base model with divided space-time attention [29]. For both TimeSformer and Video Swin Transformer, the weights are initialized from models pretrained on ImageNet-22k [31]. Each algorithm is trained using a batch size that maximizes the GPU memory capacity of two Tesla A100 GPUs. The validation set is used for evaluation after each epoch, and the epoch with the highest validation accuracy is selected for testing on an independent testing set.
For narrative generation, the VideoLLaMA2 model is employed, utilizing the CLIP ViT-Large-Patch14-336 as its video encoder and the Mistral-7B-Instruct-v0.2 as its language decoder [5]. The process begins by using the prompt “Describe this video” to extract static environment information. This information is then combined with classification results from the most effective models: SlowFast for event type classification and X-CLIP for conflict type classification. If the event type is “Normal Driving,” the narrative is generated with the prompt “Describe this event: 1: {Environment}. 2: Normal Driving.” For SCEs, the narrative is generated using the prompt “Describe this event: 1: {Environment}. 2: {Event Type}. 3: {Conflict Type}.”
4.3 Classification Task Performance
Model Learning Approach Accuracy Top5 Acc mAP AUC Balanced Acc Macro Precsion Macro F1
X-CLIP Contrastive 0.636 0.847 0.278 0.770 0.232 0.259 0.222
Action CLIP Contrastive 0.606 0.846 0.239 0.777 0.216 0.244 0.220
SlowFast Supervised 0.485 0.806 0.148 0.659 0.130 0.156 0.105
Swin Transformer Supervised 0.545 0.829 0.197 0.752 0.163 0.158 0.153
TimeSformer Supervised 0.571 0.840 0.205 0.786 0.166 0.177 0.154
Table 4: Comparison of different models in conflict type classification (Training Proportion = 0.05)
Model Learning Approach Accuracy Top5 Acc mAP AUC Balanced Acc Macro Precsion Macro F1
X-CLIP Contrastive 0.766 0.951 0.547 0.921 0.493 0.599 0.508
SlowFast Supervised 0.721 0.928 0.467 0.927 0.437 0.469 0.423
CLIP + mean pooling Supervised 0.609 0.898 0.286 0.849 0.223 0.317 0.227
CLIP + LSTM Supervised 0.611 0.872 0.243 0.843 0.227 0.212 0.215
Table 5: Comparison of alternative models in conflict type classification
Six evaluation metrics are employed to comprehensively assess the performance of the selected models. Accuracy measures the overall classification performance by using the ArgMax function to select the class with the highest predicted probability as the final output. Mean average precision (mAP) and the average area under the ROC curve (AUC) evaluate classification confidence across all thresholds. Balanced accuracy, macro precision, and macro F1 scores assess performance in imbalanced sample scenarios, emphasizing the rare-event nature of SCEs [32].
Benchmark Comparison Table 1 presents a comparative analysis of five benchmark models for four-way event type classification. All the supervised learning-based models outperformed the contrastive learning-based models in terms of accuracy, mAP, AUC, macro precision, and macro F1 score with SlowFast achieving the best performance across all evaluation metrics. This suggests that on the SHRP 2 NDS front-view video dataset, selected supervised learning approaches are more effective for event type classification task than the selected contrastive learning approaches.
Table 2 provides a comparative analysis of five benchmark models for 16-way conflict type classification. In this task, the contrastive learning-based models outperformed the supervised learning-based models in most evaluation metrics. Between the two contrastive learning-based models, X-CLIP demonstrates the best overall performance. This suggests that on the SHRP 2 NDS front-view video dataset, selected supervised learning approaches are less effective for conflict type classification than selected contrastive learning approaches.
Figure 6: Count of each conflict type with 5% training set.
Few-shot Classification Performance This experiment uses a randomly selected 5% of the training dataset to evaluate the performance of the two learning approaches in classifying conflict types under extremely limited data conditions. As shown in Figure 6, 10 out of the 16 conflict types have fewer than 10 occurrences within the 5% training dataset. By training the model on a smaller subset of data, we can assess its ability to make accurate predictions with minimal information. Table 4 provides the performance of five benchmark models for 16-way conflict type classification using 5% of the training dataset. In this limited data setting, the contrastive learning-based models significantly outperformed their supervised learning-based models across most evaluation metrics. The improvements were particularly notable in Balanced Accuracy, macro precision, and macro F1 score, indicating that the contrastive learning approach performs better on minority classes when data is scarce. This suggests that in the setting of data deficiency, selected supervised learning approaches are much less effective for conflict type classification than selected contrastive learning approaches.
Alternative model comparison To evaluate that the effective component of contrastive learning approaches in 16-way conflict type classification is the contrastive learning approach or the CLIP encoder, two alternative models are implemented. The raw video frames are processed through a CLIP image encoder, and two methods are employed to handle temporal dependencies across frames: mean pooling and long short-term memory (LSTM), similar to the methods used in ActionCLIP [24]. The video representation obtained from mean pooling or LSTM is then fed into a classifier, specifically a multi-layer perceptron, as shown in Figure 7.
Figure 7: Conflict type classification with CLIP image encoder in supervised learning approach.
Table 5 compares X-CLIP and SlowFast, the leading models for conflict type classification task using contrastive learning and supervised learning respectively. The performance of the CLIP image encoder in the supervised learning approach is notably lower, suggesting that the CLIP image encoder may not be well-suited for supervised learning approach in conflict type classification. The superior performance of the CLIP-based contrastive learning approach is largely attributed to the model’s architecture, with the text encoder playing a crucial role, especially for labels with rich text.
4.4 Prompt Tuning for VLM
Figure 8: Prompt strategies
Hallucinations in VLM refer to the scenario that response towards the prompt are fabricated, inaccurate, or misleading, often arising from the model generating responses without sufficient factual support or context [33]. To mitigate these hallucinations, approaches to design and refine prompts for VLM are proposed, involving two key steps in prompt tuning: first, instructing the VLM to describe the static environment, and second, directing the VLM to generate a narrative description of SCEs [34].
Chain-of-Thought prompt and the repetition of important information are two techniques utilized in prompt tuning [34, 35, 36]. In this study, a Chain-of Thought prompt combined with a repeat-answer strategy is employed. Initially, the VLM is prompted to describe the environment. Subsequently, the event type, conflict type, and the environment description are combined to prompt the VLM for the final narrative of the SCE, as shown in Figure 8 (c). To evaluate this approach, 20 SCEs are randomly selected from the testing set and assessed using three strategies: the proposed strategy and two alternatives. The first alternative, direct prompt (Figure 8 (a)), involves directly asking the VLM to describe the SCE based on the event type and conflict type. The second alternative, Chain-of-Thought prompt (Figure 8 (b)), involves first asking the model to describe the environment, followed by combining the event type and conflict type to generate the SCE narrative.
The narratives generated by three strategy were compared across 20 selected SCEs. The Chain-of-Thought combined with repeat-answer strategies yielded the most accurate descriptions with the fewest hallucinations. Three resulting narratives for a lead vehicle crash example on a highway are provided, with the key frames shown in Figure 9. The bold text highlights hallucinations.
Figure 9: Key frames of a lead vehicle crash example
  • Direct Prompt: The video shows a car driving on a road, and suddenly, it crashes into the lead vehicle. The collision causes the car to stop abruptly, and the driver is seen trying to regain control of the vehicle. The car then moves forward, but it is hit again by the lead vehicle, causing it to stop once again. The driver is seen trying to move the car, but it is stuck in place. The video captures the entire sequence of the crash and the driver’s attempts to recover from the situation.
  • Chain-of-Thought Prompt: The video depicts a crash and conflict with the lead vehicle. The scene unfolds on a highway where a car is seen driving behind another vehicle. Suddenly, the car collides with the lead vehicle, resulting in a crash. The impact causes the lead vehicle to veer off course, and the car that caused the crash continues to drive ahead. The conflict between the two vehicles is evident as the car that caused the crash tries to overtake the lead vehicle, which is now struggling to maintain control on the road. The video captures the intense moment of the crash and the subsequent struggle for control between the two vehicles.
  • Chain-of-Thought and repeat-answer Prompt: The video depicts a crash scenario on a busy highway. Several cars are visible, moving along the wet road, which suggests recent rain or ongoing drizzle. The environment is urban, evidenced by the presence of tall buildings in the background. The sky is overcast with clouds, adding to the dreary atmosphere. White lines are clearly painted on the road, indicating lane divisions. The crash involves a conflict with the lead vehicle, implying that one vehicle has collided with another directly in front of it. The scene is typical of a congested highway setting where traffic density and road conditions may contribute to such incidents.
Both the direct prompt and Chain-of-Thought prompt frequently reference elements not present in the video. In contrast, the Chain-of-Thought and repeat-answer prompts perform best, with the only hallucination being the mention of a wet road. This issue may arise from sunlight reflections on the road that resemble wetness. Overall, the proposed Chain-of-Thought and repeat-answer prompt strategy provides the most accurate description of the SCE.
5 Conclusion
This study introduced ScVLM, an approach that integrates supervised learning, contrastive learning, and VLM. The approach enhances the understanding of driving videos, improves the rationality of event descriptions, and reduces hallucinations in VLM-generated outputs.
Based on the SHRP 2 NDS video dataset, the results demonstrate that the proposed ScVLM generates more precise and contextually appropriate event descriptions compared to a standard VLM. This work not only contributes to the accuracy of SCE detection but also offers a robust framework for future research in automatic generation of SCE descriptions.
Acknowledgement
This project is partially funded by a grant from the National Surface Transportation Safety Center for Excellence (Grant Number: 238717).